# Judging the Judges: Evaluating Alignment and Vulnerabilities in LLMs-as-Judges

### v2

@@ Appendix B Model and dataset details
Qwen 0.5B Chat Qwen/Qwen1.5-0.5B-Chat tongyi-qianwen
Qwen 1.8B Chat Qwen/Qwen1.5-1.8B-Chat tongyi-qianwen
Qwen 4B Chat Qwen/Qwen1.5-4B-Chat tongyi-qianwen
Qwen 7B Chat Qwen/Qwen1.5-7B-Chat tongyi-qianwen
Qwen 14B Chat Qwen/Qwen1.5-14B-Chat tongyi-qianwen
Qwen 72B Chat Qwen/Qwen1.5-72B-Chat tongyi-qianwen
Appendix C Model evaluation prompt templates
@@
Appendix D Judge LLM Prompt templates
@@
Appendix E Metrics for judge models
@@
Cohen’s kappa coefficient, or Cohen’s kappa for short (Cohen, 1960), measures the alignment of two annotators while also taking into account the possibility of agreement by pure chance. This coefficient can have values between $-1$ and $1$, but is usually above 00 in most real-world situations. The value of Cohen’s kappa is given as
$\kappa$ $=\frac{p_{o}-p_{e}}{1-p_{e}}=\frac{2(T_{P}T_{N}-F_{P}F_{N})}{(T_{P}+F_{P})(T_{N}+F_{P})+(T_{P}+F_{N})(T_{N}+F_{N})}.$ (2)
Here, $p_{o}$ is the relative observed agreement, and $p_{e}$ is the hypothetical probability of chance agreement. This coefficient is considered to be a more robust measure of inter-annotator alignment, but also less interpretable in terms of what a particular value of $\kappa$ means. Generally, values of $\kappa$ in ranges $[0,0.2)$, $[0.2,0.4)$, $[0.4,0.6)$, $[0.6,0.8)$, and $[0.8,1)$ are considered to indicate no alignment, slight alignment, moderate alignment, substantial alignment, and near-perfect alignment respectively, with $\kappa=1$ indicating perfect alignment.
Appendix F Human Annotation Guidelines
@@
Appendix G Experiment costs
@@
Total 72.5 24 23
Appendix H Judge Scores
@@
Llama 3 8B 68.75 76.00 85.25 52.91 75.0 78.5 62.5 66.00 76.25
Llama 3 70B 66.00 76.50 87.00 61.52 68.50 78.25 73.5 65.00 94.50
Llama 2 7B 61.75 71.25 80.25 58.48 77.25 58.00 68.25 80.50 86.75
Llama 2 13B 53.00 61.50 66.50 46.07 27.75 39.50 57.50 39.00 73.50
Llama 2 70B 68.25 76.30 86.50 64.05 80.04 82.25 77.00 74.25 95.25
Mistral 7B 67.75 75.00 85.25 64.55 68.75 77.50 74.50 68.50 92.75
Gemma 2B 67.00 74.00 84.00 100.00 44.25 75.25 100.00 79.50 99.75
JudgeLM 52.50 60.50 64.00 41.52 32.50 60.00 57.25 45.5 68.50
GPT-4 63.25 75.00 85.0 56.7 57.50 72.0 72.5 52.75 92.25
Exact Match 46.75 56.00 63.75 24.05 0.25 36.25 59.50 20.25 58.25
Contains Match 50.75 60.00 68.00 38.98 46.25 59.50 57.25 44.00 70.00
Human Eval 62.25 72.75 83.75 56.00 56.50 72.25 71.75 60.75 91.50
Appendix I Exam-taker model base vs chat analysis
@@
Llama-2 7B 46.75 24.05 50.75 38.98 62.25 56.00 63.25 56.75 66.0 61.50
Mistral 7B 59.50 20.25 57.25 44.00 71.75 60.75 72.50 52.75 73.5 65.0
Llama-2 13B 56.00 0.25 60.00 46.25 72.75 56.50 75.00 57.50 76.50 68.50
Llama-2 70B 63.75 36.25 68.00 59.50 83.75 72.25 85.00 72.00 87.00 78.25
@@
To disentangle these two factors, we manually analyse 400 questions for Llama-2 70B and Llama-2 70B-chat, using our earlier error codes. The results, shown in Figure 10(a), sugest that, at least to some extent, the difference between base and chat models is in fact due to ‘unlearning’ of knowledge: while the number of errors is more or less equal among most categories, there is a stark difference in the incorrect entity category. Substantially more often than the base models, the chat models do answer the question with a semantically plausible but incorrect entity. In Table 7-9, we provide examples of such cases. The results do not show any evidence to support the first hypothesis: the number of errors where the answer cannot be parsed or is just entirely incorrect does not differ between base and chat models.
@@
Appendix J Exam-taker model ranking correlation
In Table 10, we show the Spearman’s rank correlation coefficient (Spearman, 1904) ($\rho$) with human judgment. Since $\rho$ > 0.7 is considered well aligned, only Llama-7B and Gemma-2B have poor rank correlation with human judgment.
@@
Judges $\rho$
Contains 0.98
JudgeLM-7B 0.98
GPT-4 0.93
Llama3-70B 0.93
Mistral-7B 0.92
Llama-13B 0.82
EM 0.78
Llama3-8B 0.77
Llama-70B 0.75
Llama-7B 0.39
Gemma-2B 0.21
Appendix K Too much info confuses judges
@@
Appendix L Judge models are sensitive to reference order
@@
Appendix M Leniency Bias
@@
The values of $P_{c}$ and $P_{+}$ can be estimated from observed data using the derived expressions. In this experiment, we include Qwen models (Bai et al., 2023) of varying sizes, in our judge ensemble to increase the number of data points for this study. The estimated probabilities using this method, with human evaluation as the reference, are shown in Figure 17(a).
To validate these derived values, we observe the correlation between the estimated values of $P_{c}$ and Cohen’s kappa ($\kappa$). As shown in Figure 17(b), we observe that the estimated values of $P_{c}$ are highly correlated to the Cohen’s kappa values for the judge models, with a Pearson correlation coefficient of $0.98$.
Judge model $\kappa$ $P_{c}$ $P_{+}$
Qwen 1.8B 0.19 0.18 1.00
Qwen 0.5B 0.19 0.19 0.87
Gemma 7B 0.40 0.32 0.07
Gemma 2B 0.50 0.62 0.80
Qwen 4B 0.69 0.67 0.89
Llama-2 7B 0.66 0.68 0.36
Mistral 7B 0.72 0.75 0.75
Llama-2 13B 0.68 0.76 0.45
Llama-2 70B 0.80 0.79 0.94
Qwen 7B 0.80 0.79 0.77
Llama-3 8B 0.81 0.80 0.84
Qwen 72B 0.82 0.80 0.93
Qwen 14B 0.83 0.82 0.83
Llama-3 70B 0.84 0.84 0.79
GPT-4 Turbo 0.85 0.85 0.66
Figure 17: a) Estimated values of $P_{c}$ and $P_{+}$ for different judge models. b) Pearson’s correlation coefficient between $\kappa$ and $P_{c}$ for judge models.

### v3

@@ Appendix C Model and dataset details
Llama-3.1 8B Chat meta-llama/Meta-Llama-3.1-8B-Instruct llama3.1
Llama-3.1 70B Chat meta-llama/Meta-Llama-3.1-70B-Instruct llama3.1
@@
Appendix D Model evaluation prompt templates
@@
Appendix E Judge LLM Prompt templates
@@
Appendix F Metrics for judge models
@@
Scott’s Pi, (Scott, 1955), measures the alignment of two annotators while also taking into account the possibility of agreement by pure chance. This coefficient usually has values above 00 in most real-world situations. The value of Scott’s Pi is given as
$\pi$ $=\frac{p_{o}-p_{e}}{1-p_{e}}=\frac{\left(\frac{TP+TN}{TP+TN+FP+FN}\right)-\left(\frac{(TP+FP)(TP+FN)+(TN+FN)(TN+FP)}{(TP+TN+FP+FN)^{2}}\right)}{1-\left(\frac{(TP+FP)(TP+FN)+(TN+FN)(TN+FP)}{(TP+TN+FP+FN)^{2}}\right)}$ (5)
Here, $p_{o}$ is the relative observed agreement, and $p_{e}$ is the hypothetical probability of chance agreement.
Appendix G Human Annotation Guidelines
@@
Appendix H Experiment costs
@@
Total 73.5 24 26
Appendix I Statistical reliability of Evaluation sample
Due to computational constraints discussed in Appendix A and Appendix H, we limit our evaluation set to randomly sampled 400 questions from TriviaQA (Joshi et al., 2017). In this section, we further take 5 samples of 300 randomly selected questions from the evaluation set and calculate the mean and standard deviation of Scott’s Pi. From Table 5, it can be observed that even on down-sampled sets, the Scott’s $\mathbf{\pi}$ values are similar to Figure 1(b). Standard deviation of all the judge models from the mean Scott’s $\mathbf{\pi}$ is also minimal, barring EM lexical match.
Table 5: Weak Scott’s $\mathbf{\pi}$ variation for the 5 down-sampled sets indicating robustness for the evaluation sample
Judge Model Mean Scott’s $\mathbf{\pi}$ Std Dev
Llama3-70B 0.88 0.0046
Llama3.1-70B 0.88 0.0039
Llama3.1-8B 0.78 0.0050
Llama2-13B 0.75 0.0043
Llama2-70B 0.69 0.0114
Mistral-7B 0.67 0.0108
JudgeLM-7B 0.66 0.0026
Contains 0.64 0.0087
Llama3-8B 0.60 0.0126
Llama2-7B 0.47 0.0112
EM 0.47 0.29
Gemma-2B 0.26 0.007
Appendix J Judge Scores
@@ Appendix I Statistical reliability of Evaluation sample
Llama 3.1 8B 65.25 75.00 83.50 60.25 70.50 75.50 73.75 59.00 89.00
Llama 3.1 70B 62.00 74.25 85.00 55.50 64.75 74.00 72.25 60.50 92.25
Llama 3 8B 76.00 83.25 91.50 73.25 82.75 85.25 81.75 76.0 97.25
Llama 3 70B 64.25 75.50 86.50 57.00 64.00 75.75 73.5 62.50 92.75
Llama 2 7B 80.50 85.25 92.00 80.50 70.75 90.75 84.00 83.25 97.75
Llama 2 13B 68.25 75.50 86.50 63.25 62.75 77.50 74.50 67.50 93.5
Llama 2 70B 71.25 80.5 90.25 67.50 74.75 81.25 80.0 72.5 96.75
Mistral 7B 72.50 80.75 90.50 69.00 74.75 82.50 80.25 72.00 96.25
Gemma 2B 79.75 87.00 91.25 58.50 41 68.50 84.0 55.75 80.50
JudgeLM 69.50 77.75 86.25 63.75 48.0 82.75 77.25 71.0 94.50
GPT-4 60.50 71.50 82.50 54.50 59.0 73.0 69.75 56.50 90.0
Exact Match 46.75 56.00 63.75 24.00 0.25 36.25 59.50 20.25 58.25
Contains Match 50.75 60.00 68.00 39.00 46.25 59.50 57.25 44.00 70.00
Human Eval 62.50 72.75 83.75 56.00 56.50 72.25 71.75 60.75 91.50
Appendix K Exam-taker model base vs chat analysis
@@
Llama-2 7B 46.75 24.00 50.75 39.00 62.25 56.00 60.50 54.50 64.25 57.00
Mistral 7B 59.50 20.25 57.25 44.00 71.75 60.75 69.75 56.50 73.50 62.50
Llama-2 13B 56.00 0.25 60.00 46.25 72.75 56.50 75.00 59.00 76.50 64.00
Llama-2 70B 63.75 36.25 68.00 59.50 83.75 72.25 82.50 73.00 86.50 75.75
@@
To disentangle these two factors, we manually analyse 400 questions for Llama-2 70B and Llama-2 70B-chat, using our earlier error codes. The results, shown in Figure 10(a), sugest that, at least to some extent, the difference between base and chat models is in fact due to ‘unlearning’ of knowledge: while the number of errors is more or less equal among most categories, there is a stark difference in the incorrect entity category. Substantially more often than the base models, the chat models do answer the question with a semantically plausible but incorrect entity. In Table 8-Table 10, we provide examples of such cases. The results do not show any evidence to support the first hypothesis: the number of errors where the answer cannot be parsed or is just entirely incorrect does not differ between base and chat models.
@@
Appendix L Exam-taker model ranking correlation
In Table 11, We use the Spearman Rank correlation coefficient (Spearman, 1904) to assess the rankings of the exam-taker models. To validate these rankings, we randomly select 6 out of 9 exam-taker models across 5 samples, subsequently calculating the mean ($\rho$) and standard deviation ($\sigma$) of the rankings. The results reveal that the contains model exhibits the highest stability and $\rho$ among the rankings, while the majority of judge models achieve a coefficient exceeding 0.7, indicating a strong alignment. Notably, smaller models such as Mistral 7B perform on par with GPT-4 Turbo, highlighting the robustness of smaller models in maintaining rankings.
@@
Judges $\rho$ $\sigma$
Contains 0.99 0.02
Mistral-7B 0.98 0.03
GPT-4 0.98 0.03
Llama2-13B 0.95 0.18
JudgeLM-7B 0.95 0.05
Llama2-7B 0.94 0.04
Llama3.1-70B 0.94 0.07
Llama3-70B 0.93 0.05
Llama3.1-8B 0.89 0.10
Llama3-8B 0.86 0.07
Llama2-70B 0.84 0.13
Gemma-2B 0.71 0.20
EM 0.67 0.13
Appendix M Too much info confuses judges
@@
Appendix N Judge models are sensitive to reference order
@@
Appendix O Leniency Bias
@@
The values of $P_{c}$ and $P_{+}$ can be estimated from observed data using the derived expressions. The estimated probabilities using this method, with human evaluation as the reference, are shown in Figure 17(a).
To validate these derived values, we observe the correlation between the estimated values of $P_{c}$ and Scott’s Pi ($\pi$). As shown in Figure 17(b), we observe that the estimated values of $P_{c}$ are highly correlated to the Scott’s $\mathbf{\pi}$ values for the judge models, with a Pearson correlation coefficient of $0.98$.
Judge model $\pi$ $P_{c}$ $P_{+}$
Gemma-2B 0.26 0.38 0.87
Llama2-7B 0.47 0.63 0.75
Llama3-8B 0.59 0.63 0.74
JudgeLM-7B 0.65 0.68 0.19
Mistral-7B 0.66 0.70 0.87
Llama2-70B 0.69 0.66 0.99
Llama2-13B 0.74 0.74 0.87
Llama3.1-8B 0.77 0.77 0.82
GPT-4 0.87 0.87 0.69
Llama3.1-70B 0.88 0.88 0.82
Llama3-70B 0.88 0.87 0.90
Figure 17: a) Estimated values of $P_{c}$ and $P_{+}$ for different judge models. b) Pearson’s correlation coefficient between $\pi$ and $P_{c}$ for judge models.